# AAAI Press Formatting Instructions for Authors Using — A Guide

### v1

@@ Introduction
The accuracy of motion is ensured by sparse control points which are proposed in SC-GS (Huang et al. 2024b). The sparse controlled points $\mathbf{P}$ and deformation network $\Phi(p_{i};\theta)$ are proposed to model the accurate deformation. Where $p_{i}$ represents the position of the $i$-th control point and $\theta$ denotes the parameters of the deformation network. The control signals at $t_{j}$ are computed by the MLP $\Phi$:
$[R_{i}^{j},T_{i}^{j}]=\Phi(p_{i},t_{j}).$ (10)
Given $K$ adjacent control points $\mathbf{p}=\{p_{0},\dots,p_{k}\}$, the 6-DoF transform $\{R^{j},T^{j}\}$ can be computed by a deformation network $\Phi$ applying on $p$, and the deformed 2D Gaussians $\mathcal{G}^{j}$ at $t_{j}$ can be interpolated using the LBS (Sumner, Schmid, and Pauly 2007) $\mathcal{I}$ including Eq. 4 and Eq. 5 :
$\mathcal{G}^{j}=\mathcal{I}\left(\{R^{j},T^{j}\},\mathcal{G}\right).$ (11)
@@
Figure 3: Mesh and mesh rendering visualization results of the D-NeRF dataset.
@@
DG-mesh* 23.474 0.903 0.103 28.045 0.972 0.088 31.667 0.977 0.036 29.217 0.965 0.049
D-3DGS 22.534 0.884 0.115 23.667 0.956 0.102 28.878 0.972 0.039 27.499 0.959 0.048
SCGS 22.784 0.882 0.113 25.231 0.956 0.103 29.534 0.976 0.035 27.095 0.961 0.048
@@
DG-mesh* 31.369 0.970 0.043 32.220 0.979 0.036 29.559 0.970 0.044 27.602 0.971 0.055
D-3DGS 28.342 0.957 0.048 29.946 0.977 0.031 28.061 0.958 0.054 25.986 0.965 0.053
SCGS 28.002 0.955 0.047 29.273 0.977 0.032 28.521 0.962 0.050 26.419 0.967 0.048
D-2DGS (Ours) 28.120 0.960 0.042 29.512 0.976 0.028 28.677 0.967 0.043 25.441 0.958 0.051
Table 1: Rendering quality of extracted meshes on the D-NeRF (Pumarola et al. 2021) dataset. The color of each cell means: best, second best. (*DG-mesh needs to query additional MLPs when rendering mesh.)
Depth filtering
During the optimization process, Geometry floaters manifest as numerous background-colored artifacts, despite the high quality of the rendered image. In the depth image, this issue is evident as noise points located outside the actual object boundaries. Consequently, after mesh extraction, a significant number of floating artifacts appear around the object.
To mitigate these artifacts, we propose a masking technique $Flitering$ as mentioned in Eq. 9 that involves extracting the object’s mask from the rendered high-quality image and applying it to filter the depth image. The process can be formalized as follows:
$\mathbf{M}(x)=\mathbf{1}[\mathbf{I}(x)=bg],$ (12)
$\mathbf{D}^{\prime}=\mathbf{D}\odot\mathbf{M},$ (13)
Loss Function
$L_{d}=\sum_{i,j}\omega_{i}\omega_{j}\left|z_{i}-z_{j}\right|,$ (14)
where $i$ indexes over intersected splats along the ray, $\omega_{i}$ denotes the blending weight of the intersection point. $\omega_{i}=\alpha_{i}\mathcal{G}_{i}(u(x)){\textstyle\prod_{j=1}^{i-1}}(1-\alpha\mathcal{G}_{j}(u(x)))$ $z_{i}$ is the depth of the intersection points.
$L_{n}=\sum_{i}\omega_{i}(1-n_{i}^{\mathrm{T}}N),$ (15)
$N(x,y)=\frac{\nabla_{x}p\times\nabla_{y}p}{\left|\nabla_{x}p\times\nabla_{y}p\right|}.$ (16)
Therefore, the overall loss function is:
$L_{all}=L_{1}(I_{r},I_{g})+\lambda_{s}L_{ssim}(I_{r},I_{g})+\lambda_{n}L_{n}+\lambda_{d}L_{d},$ (17)
Method D-NeRF datasets DG-Mesh datasets
@@
DG-Mesh* 29.114 0.963 0.057 0.698 0.123
D-3DGS 26.864 0.954 0.061 0.895 0.115
SCGS 27.107 0.954 0.060 0.856 0.115
D-2DGS(ours) 27.490 0.957 0.052 0.851 0.132
@@
Table 2: Average results on both D-NeRF datasets and DG-Mesh datasets. (*DG-mesh needs to query additional MLPs when rendering mesh.)
Mesh Extraction
The mesh at any normalized timestamp $t_{i}\in[0,1]$ can be extracted from the trained model. Given $t_{i}$, the model renders multi-view images $\mathbf{I}(t_{i})$ and depth images $\mathbf{D}(t_{i})$. The mesh is then obtained by applying TSDF using Open3D (Zhou, Park, and Koltun 2018):
$\text{Mesh}=\text{TSDF}\left(\{\mathbf{I}_{i},\mathbf{D}_{i}^{\prime}\mid i\in T\}\right).$ (18)
Experiment
Experimental Setups
We mainly compare some novel view synthesis methods for dynamic scenes based on Gaussian Splatting, including DG-Mesh (Liu, Su, and Wang 2024), Deformable 3DGS (Yang et al. 2023a), and SCGS (Huang et al. 2024b). We evaluate the quality of meshes extracted by these methods on the D-NeRF (Pumarola et al. 2021) and DG-Mesh (Liu, Su, and Wang 2024) datasets. Both datasets have large and non-rigid motions and belong to efficient multiview setups as referred to the Dycheck (Gao et al. 2022). Each timestamp has only one view which is randomly selected. Among them, the D-NeRF dataset does not contain the ground truth of the mesh but the DG-Mesh dataset includes the GT mesh. Therefore, We report the PSNR, LPIPS, and SSIM to evaluate the rendered images in the D-NeRF datasets and use Chamfer Distance (CD) and Earth Mover’s Distance (EMD) as mesh evaluation metrics. The mesh extracted by deformable 3DGS and SC-GS following the same settings as ours and 2D-GS (Huang et al. 2024a), and it is worth noting that when DG-Mesh renders an image, two additional MLPs should be maintained when querying the color of the mesh.
Figure 4: Mesh and mesh rendering visualization results of the DG-Mesh dataset.
Implementation Details
We build dynamic 2DGS upon the open-source SCGS (Huang et al. 2024b) codebase. Training iterations for all scenarios are set to 80,000. The training time will take 1 to 2 hours. The number of control points is set to 1024. All experiments are conducted on an NVIDIA RTX A5000 GPU. The hyperparameters $\lambda_{s}$, $\lambda_{n}$ and $\lambda_{d}$ are 1, 0.02 and 1000 respectively. The other hyperparameters are the same as SCGS. When the mesh is extracted, the original dynamic 2DGS representations can be removed.
Method GS render image Mesh
Color Smooth Detail
DG-Mesh
D-3DGS
SCGS
Ours
Table 3: The comparison of different models.
[…]
w/o Depth filtering 0.425 0.110
w/o Normal Consistency 0.443 0.101
w/o Depth Distortion 0.350 0.109
Full model 0.328 0.110
Table 4: Ablation experiment results of the “bird” sample in the DG-Mesh dataset.
Results
@@
The results on the D-NeRF (Pumarola et al. 2021) dataset are shown in Fig. 3. The ideal surface should be as smooth as possible while still containing the details of the object. The mesh extracted by DG-Mesh (Liu, Su, and Wang 2024) has a low resolution and lacks some details of objects, such as the hands of the ‘jumpingjacks’. Deformable-3DGS and SCGS both use 3D Gaussians as the basic representational primitives, which lack geometric accuracy. The extracted mesh overfits the details and is not geometrically accurate and smooth, such as the heads of the ‘jumpingjacks’. The three comparative methods are unable to simulate smooth planar and spherical surfaces, as demonstrated by the “bouncingballs” example. The sphere appears broken and the chassis is uneven due to the influence of lighting. In comparison, our Dynamic 2DGS highlights reconstructing rich details and modeling smooth surfaces.
@@
The rendering results of the mesh are shown in Tab. 1. When DG-Mesh renders an image, it needs to query additional MLPs to obtain the color of the mesh, thus achieving higher PSNR and SSIM. The proposed D-2DGS achieved higher indicators compared with D-3DGS and SCGS.
@@
As shown in Fig. 4, on the DG-Mesh dataset, the proposed method can extract smoother meshes compared to D-3DGS and SCGS. At the same time, compared to DG-Mesh, the proposed method can show the details of objects, such as the mouth of the ‘horse’. The average metrics of the D-NeRF and DG-Mesh datasets are shown in Tab. 2. The data in the DG-Mesh dataset has less details, so the proposed model is comparable to other models in terms of metrics.
We summarize the methods on the Tab. 3. DG-Mesh produces low-quality rendered images through Gaussian rendering, and the extracted meshes rely on MLP modeling for color, and lacking in detail. Meanwhile, D-3DGS and SCGS extract meshes that are over-fitted to details, resulting in surfaces that are not sufficiently smooth.
Figure 5: Ablation Study on Sparse point control.
Ablation Study
Sparse point control.
We conducted ablation experiments on sparse point control, where we removed the sparse point control and directly modeled the offset for 2DGS using an MLP as the offset field. As shown in Fig. 5, in areas with significant motion, such as a person’s hand and a horse’s hoof, sparse point control allows for the extraction of a more detailed mesh.
Figure 6: Ablation Study on Depth filtering.
Depth filtering.
We removed the mask on the Depth image and then conducted experiments. As can be seen in Fig. 6 (a), there are many noise points around the object in the rendered depth image, which may be caused by the offset field offsetting the 2D Gaussians. After extracting the mesh through TSDF fusion, some floating objects will appear around the mesh of the object. With the help of high-quality renderings, the object mask is extracted and the depth image is masked. The noise in the depth image can be filtered out, thereby removing the floating objects in the mesh, as shown in Fig. 6 (b).
Loss function.
We conduct ablation experiments on various loss functions to verify the effectiveness of the loss functions. As shown in Fig. 7, the two regularization terms in the loss function are important for extracting high-quality meshes. The Normal Consistency regularization term can make the normals of Gaussians more accurate, and the Depth Distortion term can make the Gaussians distributed on the plane. The corresponding metrics are shown in Tab. 4.
Limitations and Future work
Though dynamic 2D Gaussians can model the accurate dynamic surface, there are several limitations: (1) Some holes exist in the extracted mesh. (2) Given the extremely sparse views, such as strict monocular input, robustness priors can other low-rank motion representations (Wang et al. 2024a) could be explored to model more accurate motion. (3) More post-processing methods could be extended to handle Geometry floaters and holes completion.
Figure 7: Ablation Study on loss function.
Conclusion
In this paper, we propose Dynamic 2D Gaussians to extract high-quality meshes from dynamic 2D images input. We use 2D Gaussians as the basic representation primitives and use sparse control points to control 2D Gaussians to model the motion of objects. We used high-quality rendered images to filter out the noise in the depth image and used TSDF to obtain higher-quality meshes. Experiments show that the mesh extracted by the proposed method has high details and a smoother surface.
Appendix
Written by AAAI Press Staff1With help from the AAAI Publications Committee.
AAAI Style Contributions by Pater Patel Schneider, Sunil Issar,
[…]
Figure 8: Relighting and composition results in generated mesh on the Blender.
Figure 9: Rendered images, normals, and depth maps.
Figure 10: Comparison of results with different background colors.
In the supplementary material, we mainly provide additional experimental results in Sec. 1. Then more discussions are conducted in Sec. 2. Finally, we delve into the limitations of our proposed Dynamic 2D Gaussians in Sec. 3.
Additional Experimental Results
In the D-NeRF dataset, the metrics of the RGB images rendered from the extracted meshes do not fully reflect the quality of the meshes. Therefore, we provide additional visualization results, as shown in Fig. 11. DG-Mesh exhibits relatively poorer performance in modeling object details. The meshes extracted by Deformable-3DGS and SCGS exhibit numerous pits and holes on the surface. In comparison, the meshes extracted by the proposed Dynamic 2D Gaussians approach capture object details while maintaining a smoother surface, reflecting better geometric properties.
Figure 11: Mesh and mesh rendering visualization results of the D-NeRF dataset.
In the DG-Mesh dataset, the detailed metrics are shown in Table 1. Since the data of the DGMesh dataset is relatively simple and the surface is smooth, it is difficult to reflect the advantages of the proposed method in terms of metrics. Our metrics are close to those of SCGS, but the visual quality shows that our extracted surfaces appear smoother than SCGS. The DG-Mesh method shows advantages in the Duck and Torus2sphere examples, but performs relatively poorly on the other cases. Although DGMesh has a high metrics in Torus2sphere, from the visualization results in Fig. 12, DGMesh does not show a particularly good reconstruction effect in the Torus2sphere example.
The proposed method can render Gaussians to obtain normal depth maps and high-quality RGB images of dynamic objects, as shown in Fig. 9.
@@ Conclusion
D-3DGS 1.064 0.076 0.298 0.114 0.427 0.123
SCGS 1.001 0.074 0.232 0.118 0.336 0.121
DG-mesh 0.782 0.048 0.299 0.168 0.557 0.128
Ours 1.040 0.092 0.391 0.177 0.328 0.110
@@
D-3DGS 0.585 0.101 2.655 0.155 0.340 0.122
SCGS 0.528 0.100 2.716 0.155 0.321 0.124
DG-mesh 0.626 0.114 1.515 0.126 0.406 0.153
Ours 0.544 0.122 2.479 0.164 0.324 0.129
@@
Table 5: Mesh reconstruction quality comparison results of DG-Mesh dataset.
Figure 12: Visualization results of DGMesh dataset.
Discussion and Application
Background Color.
Training models with different background colors can have some influence on the extraction of the mesh. The background color should be selected to have significant contrast with the color of the object itself. For the training images, the background color should be selected to have significant contrast with the color of the object itself. As shown in Fig. 10, if a black background is used, the shadows on the object are difficult to distinguish from the background, making it challenging to extract the object’s mask from the rendered images. This then makes it difficult to filter out floating artifacts in the depth maps, ultimately leading to the presence of many floating artifacts in the extracted meshes.
Mesh composition and relighting.
The mesh reconstructed by Dynamic 2D Gaussians can be applied in the Blender in the Fig. 8. Placing a proper light source, the shadow exists in the image due to the occlusion relationship between the objects.
Time-Aware Mesh.
Our method supports the extraction of dynamic mesh sequences, as shown in Fig. 13. The extracted precise temporal geometry provides the potential for analyzing the motion characteristics and dynamics of objects.
Figure 13: Time-Aware Mesh.
Limitations
Due to the thin nature of 2DGS and the sparsity of dynamic scene views, the extracted mesh may be broken in some areas. As shown in Fig. 14, the mesh is prone to damage in the following two situations: 1. In some places that are difficult to observe from the training perspective and in shadows, the mesh is prone to damage, such as Fig. 14 (a). 2. Mesh is prone to damage on uneven thin planes with uneven lighting, such as Fig. 14 (b). Subsequent research is needed on better methods for representing precise dynamic geometry, as well as targeted mesh damage repair techniques. This will help us better extract mesh from dynamic objects.
Figure 14: Some areas of the mesh will appear broken.

### v2

@@ 4.2. Sparse point controlled 2D Gaussians
The accuracy of motion is ensured by sparse control points, which are proposed in SC-GS (Huang et al., 2024a). The sparse controlled points $\mathbf{P}$ and deformation network $\Phi(p_{i};\theta)$ are proposed to model the accurate deformation. Where $p_{i}$ represents the position of the $i$-th control point and $\theta$ denotes the parameters of the deformation network. The control signals at $t_{j}$ are computed by the MLP $\Phi$:
(10) $R_{i}^{j},T_{i}^{j}]=\Phi(p_{i},t_{j}).$
Given $K$ adjacent control points $\mathbf{p}=\{p_{0},\dots,p_{k}\}$, the 6-DoF transform $\{R^{j},T^{j}\}$ can be computed by a deformation network $\Phi$ applying on $p$, and the deformed 2D Gaussians $\mathcal{G}^{j}$ at $t_{j}$ can be interpolated using the LBS (Sumner et al., 2007) $\mathcal{I}$ including Eq. 4 and Eq. 5 :
(11) $\mathcal{G}^{j}=\mathcal{I}\left(\{R^{j},T^{j}\},\mathcal{G}\right).$
@@
Figure 4. Mesh and mesh rendering visualization results of the D-NeRF dataset.
4.3. Depth filtering
During the optimization process, Geometry floaters manifest as numerous background-colored artifacts, despite the high quality of the rendered image. In the depth image, this issue is evident as noise points located outside the actual object boundaries. Consequently, after mesh extraction, a significant number of floating artifacts appear around the object. To mitigate these artifacts, we propose a masking technique $Flitering$ as mentioned in Eq. 9 that involves extracting the object’s mask from the rendered high-quality image and applying it to filter the depth image. The process can be formalized as follows:
(12) $\mathbf{M}(x)=\mathbf{1}[\mathbf{I}(x)=bg],$
(13) $\mathbf{D}^{\prime}=\mathbf{D}\odot\mathbf{M},$
4.4. Loss Function
(14) $L_{d}=\sum_{i,j}\omega_{i}\omega_{j}\left|z_{i}-z_{j}\right|,$
where $i$ indexes over intersected splats along the ray, $\omega_{i}$ denotes the blending weight of the intersection point. $\omega_{i}=\alpha_{i}\mathcal{G}_{i}(u(x)){\textstyle\prod_{j=1}^{i-1}}(1-\alpha\mathcal{G}_{j}(u(x)))$ . $z_{i}$ is the depth of the intersection points.
(15) $L_{n}=\sum_{i}\omega_{i}(1-n_{i}^{\mathrm{T}}N),$
(16) $N(x,y)=\frac{\nabla_{x}p\times\nabla_{y}p}{\left|\nabla_{x}p\times\nabla_{y}p\right|}.$
Therefore, the overall loss function is:
(17) $L_{all}=L_{1}(I_{r},I_{g})+\lambda_{s}L_{ssim}(I_{r},I_{g})+\lambda_{n}L_{n}+\lambda_{d}L_{d},$
4.5. Mesh Extraction
The mesh at any normalized timestamp $t\in[0,1]$ can be extracted from the trained model. Given $t$, the model renders multi-view images $\mathbf{I}(t)$ and filtered depth images $\mathbf{D}^{\prime}(t)$. The mesh is then obtained by applying TSDF (Zhou et al., 2018):
(18) $\text{Mesh}(t)=\text{TSDF}\left(\mathbf{I}(t),\mathbf{D}^{\prime}(t)\right).$
The mesh of an object at any time can be extracted by changing the $t$ value.
4.6. Post-processing
Experiments have found that the extracted mesh occasionally has problems with damage and holes, so we use MeshLab to repair small holes in the extracted mesh. The hole-filling algorithm first detects the holes on the mesh, then sets the maximum hole size allowed to be filled, and finally generates new triangular patches to close the holes for the hole boundaries that meet the conditions.
5. Experiment
5.1. Experimental Setups
We mainly compare some novel view synthesis methods for dynamic scenes, including TiNeuVox-B (Fang et al., 2022), DGMesh (Liu et al., 2025), DynaSurfGS (Cai et al., 2024), Deformable 3DGS (Yang et al., 2024a), and SCGS (Huang et al., 2024a). We evaluate the quality of meshes extracted by these methods on the D-NeRF (Pumarola et al., 2021) and DGMesh (Liu et al., 2025) datasets. Both datasets have large and non-rigid motions and belong to efficient multiview setups as referred to the Dycheck (Gao et al., 2022). Each timestamp has only one view, which is randomly selected. Among them, the D-NeRF dataset does not contain the ground truth of the mesh. Therefore, we report the PSNR, LPIPS, and SSIM of the rendered images by mesh to evaluate the mesh quality in the D-NeRF datasets. For the DGMesh dataset, we report the CD and EMD to evaluate the quality of the generated mesh. It is worth noting that when DGMesh renders an image, two additional MLPs should be maintained when querying the color of the mesh. In addition, DGMesh uses a differentiable Poisson Solver and differentiable Marching Cubes to extract the mesh. Other methods use TSDF to extract mesh like 2D-GS (Huang et al., 2024b).
5.2. Implementation Details
We build dynamic 2DGS upon the open-source SCGS (Huang et al., 2024a) codebase. Training iterations for all scenarios are set to 80,000. The training time will take 1 to 2 hours. The number of control points is set to 1024. All experiments are conducted on an NVIDIA RTX A5000 GPU. The hyperparameters $\lambda_{s}$, $\lambda_{n}$ and $\lambda_{d}$ are 1, 0.02 and 1000 respectively. The other hyperparameters are the same as SCGS. When the mesh is extracted, the original dynamic 2DGS representations can be removed. The indicators of TiNeuVox-B and DGMesh come from the DGMesh paper.
@@ 5.2. Implementation Details
TiNeuVox-B (Fang et al., 2022) 21.927 0.843 0.126 24.819 0.947 0.101 23.621 0.932 0.075 21.429 0.908 0.085
D-3DGS (Yang et al., 2024a) 22.534 0.884 0.115 23.667 0.956 0.102 28.878 0.972 0.039 27.499 0.959 0.048
SCGS (Huang et al., 2024a) 22.784 0.882 0.113 25.231 0.956 0.103 29.534 0.976 0.035 27.095 0.961 0.048
DGMesh* (Liu et al., 2025) 21.289 0.838 0.159 29.145 0.972 0.099 31.769 0.977 0.045 27.884 0.954 0.074
@@
TiNeuVox-B (Fang et al., 2022) 22.967 0.925 0.064 24.263 0.941 0.051 24.219 0.927 0.070 18.657 0.917 0.118
D-3DGS (Yang et al., 2024a) 28.342 0.957 0.048 29.946 0.977 0.031 28.061 0.958 0.054 25.986 0.965 0.053
SCGS (Huang et al., 2024a) 28.002 0.955 0.047 29.273 0.977 0.032 28.521 0.962 0.050 26.419 0.967 0.048
DGMesh* (Liu et al., 2025) 30.400 0.968 0.055 30.208 0.974 0.051 28.951 0.959 0.065 25.460 0.959 0.084
D-2DGS (Ours) 28.120 0.960 0.042 29.512 0.977 0.028 28.677 0.967 0.043 25.498 0.958 0.049
Table 1. Rendering quality of extracted meshes on the D-NeRF (Pumarola et al., 2021) dataset (Background: white). The color of each cell means: best, second best. (*DGMesh uses differentiable Marching Cubes to extract mesh, and the extracted mesh lacks details.)
Figure 5. Mesh visualization results of the DGMesh dataset.
Method D-NeRF datasets DGMesh datasets
@@
TiNeuVox-B (Fang et al., 2022) 22.738 0.918 0.086 1.839 0.130
D-3DGS (Yang et al., 2024a) 26.864 0.954 0.061 0.895 0.115
SCGS (Huang et al., 2024a) 27.107 0.954 0.060 0.856 0.115
DGMesh* (Liu et al., 2025) 28.137 0.950 0.057 0.698 0.123
DynaSurfGS (Cai et al., 2024) - 0.910 0.123
D-2DGS(Ours) 27.498 0.957 0.052 0.647 0.112
@@
Table 2. Average results on both D-NeRF datasets and DGMesh dataset.
5.3. Results
@@
The results on the D-NeRF (Pumarola et al., 2021) dataset are shown in Fig. 4. The ideal surface should be as smooth as possible while still containing the details of the object. The mesh extracted by DGMesh (Liu et al., 2025) has a low resolution and lacks some details of objects, such as the hands of the ‘jumpingjacks’. Deformable-3DGS and SCGS both use 3D Gaussians as the basic representational primitives, which lack geometric accuracy. The extracted mesh overfits the details and is not geometrically accurate and smooth, such as the heads of the ‘jumpingjacks’. The three comparative methods are unable to simulate smooth planar and spherical surfaces, as demonstrated by the “bouncingballs” example. The sphere appears broken, and the chassis is uneven due to the influence of lighting. In comparison, our Dynamic 2DGS highlights reconstructing rich details and modeling smooth surfaces.
@@
The rendering results of the mesh are shown in Tab. 1. When DGMesh renders an image, it needs to query additional MLPs to obtain the color of the mesh, thus achieving higher PSNR. The proposed D-2DGS achieved higher indicators. The average metrics are shown in Tab. 2 and Tab. 4.
@@
As shown in Fig. 5, on the DGMesh dataset, the proposed method can extract smoother meshes compared to D-3DGS and SCGS. Compared to DGMesh, the proposed method can show the details of objects, such as the mouth of the ‘horse’. The indicators on the DGMesh dataset are shown in Tab. 3 and Tab. 2. The proposed method achieves the lowest CD and EMD value, demonstrating the effectiveness of the proposed method on this dataset.
We evaluate the quality of the mesh from three aspects: whether it is clean (i.e., whether there are floating objects), whether it is smooth, and whether it has details. The image quality rendered by the DGMesh method through Gaussian is relatively low. In addition, the mesh extracted by DGMesh lacks details. The mesh extracted by D-3DGS and SCGS is prone to floaters. In addition, its surface is uneven, prone to potholes, and not smooth enough. In comparison, the mesh extracted by our method not only has better details but is also relatively smooth.
@@
TiNeuVox-B (Fang et al., 2022) 0.969 0.059 1.918 0.246 8.264 0.215
D-3DGS (Yang et al., 2024a) 1.064 0.076 0.298 0.114 0.427 0.123
SCGS (Huang et al., 2024a) 1.001 0.074 0.232 0.118 0.336 0.121
DGMesh (Liu et al., 2025) 0.782 0.048 0.299 0.168 0.557 0.128
D-2DGS (Ours) 0.841 0.057 0.327 0.134 0.328 0.110
@@
TiNeuVox-B (Fang et al., 2022) 0.874 0.129 2.115 0.203 0.568 0.184
D-3DGS (Yang et al., 2024a) 0.585 0.101 2.655 0.155 0.340 0.122
SCGS (Huang et al., 2024a) 0.528 0.100 2.716 0.155 0.321 0.124
DGMesh (Liu et al., 2025) 0.626 0.114 1.515 0.126 0.406 0.153
D-2DGS (Ours) 0.553 0.115 1.502 0.139 0.332 0.116
@@
Table 3. Mesh reconstruction quality comparison results of DGMesh dataset.
Method Train Time Render Speed Memory Storage
TiNeuVox-B (Fang et al., 2022) 28 mins 1.5 FPS 5574 MB 48 MB
D-3DGS (Yang et al., 2024a) 10 mins 190 FPS 4270 MB 10 MB
SCGS (Huang et al., 2024a) 16 mins 215 FPS 3200 MB 26 MB
DGMesh (Liu et al., 2025) 90 mins 140 FPS 9425 MB 17 MB
D-2DGS (Ours) 32 mins 71 FPS 3690 MB 20 MB
Table 4. Comprehensive comparison of different methods.
Figure 6. Ablation Study on Sparse point control.
5.4. Ablation Study
Sparse point control.
We conducted ablation experiments on sparse point control, where we removed the sparse point control and directly modeled the offset for 2DGS using an MLP as the offset field. As shown in Fig. 6, in areas with significant motion, such as a person’s hand and a horse’s hoof, sparse point control allows for the extraction of a more detailed mesh. This shows that sparse point control can model finer motions, thus helping to extract more detailed meshes.
Figure 7. Ablation Study on Depth filtering.
Depth filtering.
We removed the mask on the Depth image and then conducted experiments. As can be seen in Fig. 7 (a), there are many noise points around the object in the rendered depth image, which may be caused by the offset field offsetting the 2D Gaussians. After extracting the mesh through TSDF fusion, some floating objects will appear around the mesh of the object. With the help of high-quality renderings, the object mask is extracted, and the depth image is masked. The noise in the depth image can be filtered out, thereby removing the floaters in the mesh, as shown in Fig. 7 (b). Experiments show that the proposed depth filtering method is beneficial for removing floaters and obtaining a cleaner mesh.
Loss function.
We conduct ablation experiments on various loss functions to verify the effectiveness of the loss functions. As shown in Fig. 8, the two regularization terms in the loss function are important for extracting high-quality meshes. Among them, the normal consistency regularization term can make the normal of the Gaussian sphere more accurate, thereby making the surface of the mesh smoother. The Depth Distortion term can make the Gaussians distributed on the plane, thus reducing the surface damage. The corresponding metrics are shown in Tab. 5. It can be found from the table that adding two regularization terms to the loss function significantly reduces the CD value and improves the quality of the mesh.
Figure 8. Ablation Study on loss function.
Post-processing.
Since the input view of the model is sparse, the extracted mesh is prone to holes in some areas that are difficult to see. Therefore, we use some post-processing methods to repair the holes that may appear in the mesh to improve the integrity of the mesh. As shown in Fig. 9, we used MeshLab to repair the holes in the mesh. As can be seen from the figure, simple post-processing operations can fill in the holes that may appear in the reconstructed mesh.
Figure 9. Repair holes in the mesh.
$downarrow Method CD\downarrow EMD\downarrow$
w/o DF 0.425 0.110 w/o DD 0.350 0.109
w/o NC 0.443 0.101 Full 0.328 0.110
Table 5. Ablation experiment results of the “bird” sample in the DGMesh dataset. (DF: Depth filtering, NC: Normal Consistency, $D$: Depth Distortion)
6. Limitations and Future work
Though dynamic 2D Gaussians can model the accurate dynamic surface, there are several limitations: (1) Some holes exist in the extracted mesh. We need some post-processing operations to repair these holes. Future work will pursue direct hole-free generation to bypass repair entirely. (2) Under ultra-sparse (e.g., monocular) views, robust priors and low-rank motion can enhance accuracy. Recently, 3D foundation models (Xiang et al., 2025; Tochilkin et al., 2024; Li et al., 2025; Ye et al., 2025) show the great potential to generate high quality mesh from a static image. We believe that reconstruct high quality mesh from the dynamic multiview input will be coming soon.
7. Conclusion
In this paper, we propose Dynamic 2D Gaussians to extract high-quality meshes from dynamic 2D images input. We use 2D Gaussians as the basic representation primitives and use sparse control points to control 2D Gaussians to model the motion of objects. We extract masks from high-quality rendering data and use these masks to filter out noise in the depth map. Finally, we use TSDF to extract the mesh of the object at any time. Experiments show that the mesh extracted by the proposed method has high details and a smoother surface.
Acknowledgements.
This work was supported by National Natural Science Foundation of China (No. 62376102).